Calculate the emf of a Cd-Cu cell in which Cd is in contact with 0.002 M CdSO4 and Cu in contact with 0.02 M CuSO4 solution. The standard emf of the cell is 0.74 V at 298 K.
1
Expert's answer
2019-09-10T03:47:32-0400

The cell reaction is

"Cd +Cu^{+2} \to Cd^{+2}+Cu"

Number of electrons used"=2"

"E_{cell}=E_0-\frac{0.0591}{n} \space log_{10} \frac{[Cd^{+2}][Cu]}{[Cu^+][Cd]}"

"[Cd]=[Cu]=1"

"E_{cell}=0.74-\frac{0.0591}{2} \space log_{10} \frac{[Cd^{+2}]}{[Cu^{+2}]}"

"E_{cell}=0.74-\frac{0.0591}{2} \space log_{10} \frac{0.002}{0.02}=0.74+0.0295=0.7695 V"